OpenStudy (anonymous):

What is the sum of the finite arithmetic series? (–10) + 0 + 10 + 20 + … + 130

OpenStudy (tkhunny):

You should have a formula for that. Without such a formula, is your algebra up tp speed so you can CREATE it? -10 = -10 + 0 0 = -10 + 10 10 = -10 + 2*10 20 = -10 + 3*10 ... 130 = -10 + 14*10 That's 15 terms, so -10*15 + 10(0 + 1 + 2 + 3 + ... + 14) = -10*15 + 10(15*7) = 10*15*(-1 + 7) Did we get anywhere?

OpenStudy (anonymous):

Sum of {-10+(10*k)} from k=0 to infinity.

OpenStudy (anonymous):

Do you mean this formula, I just learned it \[\large s_n = \dfrac{n(a_1+a_n)}{2}\] \[\large s_{15} = \dfrac{15(-10+130)}{2}\] \[\large s_{15} = \dfrac{15(-10+130)}{2}\] \[\large s_{15} = 900\]

OpenStudy (anonymous):

Formula is {C+dK} from K=0 to infinity C = first term. -10 d = 10 K = sequence #

OpenStudy (tkhunny):

Will you look at that?! 6*10*15 = 900. Almost magic!

OpenStudy (anonymous):

Was that correct? Please medal if it was correct.

OpenStudy (tkhunny):

@1Rascal Good try, but an infinite arithmetic series with a difference of 10 will not be adding up to anything anytime soon.